# New results on $k$-Roudenff’s conjecture

### v1

@@ 2.1. $k$-neighborly Lawrence oriented matroid (LOM)
A Lawrence oriented matroid (LOM) of rank $r$ on the totally ordered set $E=\{1,\ldots,n\}$, $r\leq n$, is an uniform oriented matroid obtained as the union of $r$ uniform oriented matroids ${\mathcal{M}}_{1},\ldots,{\mathcal{M}}_{r}$ of rank $1$ on $(E,<)$ (see [17, 18]). They were studied and used in [32] to find examples to provide upper bounds to the McMullen problem, stated in [23], and then in a generalization of this problem to $k$-neighborly polytopes [14]. Note that yet another variant of McMullen’s problem has been studied recently in which LOMs are also used [15]. In [29] they were also used to prove Roudneff’s conjecture (Conjecture 1.1) for LOMs.
@@
First, notice that the alternating oriented matroid is a LOM, ${C}_{r}(n)=\mathcal{M}_{A}$, where $A$ is the matrix with all its entries of the same sign (in fact, we point out in Subsection 2.4 that $A$ is not unique). Second, acyclic LOMs are realizable since they are unions of realizable oriented matroids [3, Proposition 8.2.7]. Thirdly, as well as the signs of the bases of a LOM $\mathcal{M}_{A}$ can be obtained directly from matrix $A$ (Equation (2)), it is also possible to obtain from $A$ the signs of the elements of any circuit of $\mathcal{M}_{A}$ in the following way. It is well-known that the chirotope of a uniform oriented matroid relates the basis and the circuits as follows: $\chi(B)=-X_{j_{i}}\cdot X_{j_{i+1}}\cdot\chi(B^{\prime}),$ where $\underline{X}=\{j_{1},...,j_{r+1}\}$, $B=\underline{X}\setminus j_{i}$ and $B^{\prime}=\underline{X}\setminus j_{i+1}$ see [3, Section 3.5]. Thus, it can be seen that the signs of any circuit $\underline{X}=\{j_{1},...,j_{r+1}\}$ in ${\mathcal{M}}_{A}$ can be obtained from matrix $A$ by the equation
@@
see [32, Lemma 2.1]. We also note the following.
@@
The Top and Bottom travels of a matrix $A$ were defined and studied in [32] to detect when $\mathcal{M}_{A}$ is $1$-neighborly. Next, we will use them to determine when $\mathcal{M}_{A}$ is $k$-neighborly (Lemma 2.3).
@@
[32, Lemma 3.1] There is a bijection between the set of all plain travels of the matrix $A$ and the set of all acyclic reorientations of $\mathcal{M}_{A}$.
@@
In this subsection we answer Conjecture 1.2 in the affirmative for LOMs when $r=2k+2\geq 4$ and $n\geq 3r-4$ (Theorem 2.11). The case $r=2k+1\geq 3$ and $n\geq r+2$ was proved for any oriented matroid in [21, Corollay 3.17].
Let denote by $PT$ the set of plain travels of a matrix $A$ and let $PT^{k}$ be the set of plain travels of $A$ that are $k$-neighborly. More generally, given any set $\mathcal{P}$ of plain travels of $\mathcal{M}$, we denote by $\mathcal{P}^{k}$ the subset of plain travels of $\mathcal{P}$ that are $k$-neighborly. We notice that for a LOM $\mathcal{M}$, we have that $f_{\mathcal{M}}(r,n,k)=2|PT^{k}|$ by Lemma 2.5 and Remark 2.6. Thus, we may rewrite [21, Corollay 3.17] now in terms of $k$-neighborly plain travels as follows.
@@ 2.5. $k$-Roudneff’s conjecture for low ranks
In this section we prove $k$-Roudneff’s Conjecture for LOMs for the cases $r=7$, $k=2$ and $n\geq 11$ (Corollary 2.26), for $r=8$, $k=2,3$ and $n=11,12$ (Theorem 2.27), for $r=9$, $k=2$ and $n=12$ and also for $r=9$, $k=3$ and $n\geq 12$ (Corollary 2.29). We recall that 1.2 was solved for LOMs when $k=1$ in [29].
In [21] it was proved that 1.2 can be reduced to a finite number of cases as shown in the following theorem.
@@
[21, Theorem 3.16] If $f_{\mathcal{M}^{\prime}}(r^{\prime},n^{\prime},k)\leq c_{r^{\prime}}(n^{\prime},k)$ for every rank $r^{\prime}\leq r$ uniform oriented matroid $\mathcal{M}^{\prime}$ on $n^{\prime}=2(r^{\prime}-k)+1$ elements, then
@@
[21, Lemma 3.15] Let $\mathcal{M}$ be rank $r\geq 2k+1\geq 3$ uniform oriented matroid and let $e\in E$. Then,
@@
[3, Proposition 7.6.7] The family of LOMs is closed under contraction and deletion of elements.
@@
In [21, Subsection 4.1], the authors provided a computer program to obtain $f_{\mathcal{M}}(r,n,k)$ from its chirotope, available at [1] (in fact they obtained it via something called the $o$-vector whose $i$-th entry is $o_{i}$ with $i=0,\ldots,\lfloor\frac{r-1}{2}\rfloor$ and $o_{i}$ is the number of $i$-neighborly reorientations of $\mathcal{M}$ that are not $i+1$-neighborly). Now, for a LOM $\mathcal{M}_{A}$, we may use Equation (2) to obtain the signature of its basis an so its chirotope. Thus, using the computer program mentioned above we obtain $f_{\mathcal{M}_{A}}(r,n,k)$.
By Remark 1.3 it is enough to check Conjecture 1.2 only for the reorientation classes $\mathcal{M}$] of $\mathcal{M}$. So, in the case of LOMs it is sufficient by Theorem 2.18 to choose a representative matrix of a chessboard to obtain a representative matrix of its corresponding reorientation class. Hence, the computer program (available at [1]) has to compute $2^{(n-r-1)(r-1)}$ different chirotopes to obtain $f_{\mathcal{M}_{A}}(r,n,k)$ for the case of LOMs by Theorem 2.20.
@@ 2.5.2. $k$-Roudneff’s conjecture for LOMs with $r=7$, $k=2$ and $n\geq 11$
We first notice that 1.2 was already proved for $r=7$, $k=2$ and $n=10$ in [21, Theorem 4.1 (d)]. Now, in order to prove the conjecture for LOMs with $r=7$, $k=2$ and $n\geq 11$ it is enough to prove only the case $n=11$ by Proposition 2.24 since for $r<7$ and $k=2$ 1.2 was proved in [21]. As mentioned above, we have to analyze $2^{(n-r-1)(r-1)}=2^{18}=262$ $144$ different matrices. We implemented our computer program and we resume the results in the following theorem.
@@
On the other hand, notice that 1.2 holds for $r=7$, $k=3$ since it was already proved for any oriented matroid in [21, Theorem 3.17].
@@ 2.5.3. $k$-Roudneff’s conjecture for LOMs with $r=8$, $k=2,3$ and $n=11,12$
On the other hand, notice that 1.2 holds for $r=9,k=4$ since it was already proved for any oriented matroid in [21, Theorem 3.17]. Finally, we observe that in order to prove the conjecture for LOMs with $r=9$ $k=2$ and $n\geq 15$ (using Proposition 2.24) the computer program has to calculate $2^{40}$ different matrices by Theorem 2.20.
@@ 3. $k$-Roudneff’s conjecture holds asymptotically
We will prove in Theorem 3.9 that $f(r,n,k)\leq F(r,n,k)$ for every $r\geq 2k+2\geq 4$ and $n\geq r+1\geq 5$ but first, we need some previous results. $k$-Roudnedff’s conjecture was proved for $r=2k+1\geq 3$ and $n\geq r+2$ in [21, Corollary 3.17]. On the other hand, the value of $c_{r}(n,k)$ was computed for $n\geq r+1$ when $r=2k+1$ in [21, Theorem 3.11 and Proposition 3.12]. Combining these results and Remark 1.3 we obtain the following result that will be useful in this section.
@@
[21] Let $\mathcal{M}$ be a rank $r=2k+1\geq 3$ oriented matroid on $n\geq r+1$ elements, then
@@
The value of $f(r,r,k)$ makes sense if we consider the oriented matroid as an arrangement. As we pointed out in the introduction (Subsection 1.2), an oriented matroid on $n$ elements of rank $r$ is naturally associated with some arrangement of $n$ (pseudo) hyperplanes in the projective space of dimension d=r1d=r-1. The arrangement decomposes the projective space into $d$-cells which in turn correspond to the acyclic reorientations (i.e., 0-neighborly reorientations) of the oriented matroid (We refer the reader to [11] to more details on the Topological Representation Theorem). In this sense, it is well-known that $f(r,r,0)\leq 2^{r-1}$ (see [5]).
@@ Acknowledgements.
R. Hernández-Ortiz and L. P. Montejano were supported by SGR Grant 2021 00115 and by the project HERMES, funded by INCIBE and the European Union NextGenerationEU/PRTR. L. P Montejano was supported by the Ministry of Education and Science, Spain, under the project PID2023-146925OB-I00.

### v2

@@ 2.1. $k$-neighborly Lawrence oriented matroid (LOM)
A Lawrence oriented matroid (LOM) of rank $r$ on the totally ordered set $E=\{1,\ldots,n\}$, $r\leq n$, is an uniform oriented matroid obtained as the union of $r$ uniform oriented matroids ${\mathcal{M}}_{1},\ldots,{\mathcal{M}}_{r}$ of rank $1$ on $(E,<)$ (see [lawrence1981unions, RS88]). They were studied and used in [R01] to find examples to provide upper bounds to the McMullen problem, stated in [L72], and then in a generalization of this problem to $k$-neighborly polytopes [GL15]. Note that yet another variant of McMullen’s problem has been studied recently in which LOMs are also used [garcia2023number]. In [MR15] they were also used to prove Roudneff’s conjecture (Conjecture 1.1) for LOMs.
@@
First, notice that the alternating oriented matroid is a LOM, ${C}_{r}(n)=\mathcal{M}_{A}$, where $A$ is the matrix with all its entries of the same sign (in fact, we point out in Subsection 2.4 that $A$ is not unique). Second, acyclic LOMs are realizable since they are unions of realizable oriented matroids [BVSWZ99, Proposition 8.2.7]. Thirdly, as well as the signs of the bases of a LOM $\mathcal{M}_{A}$ can be obtained directly from matrix $A$ (Equation (2)), it is also possible to obtain from $A$ the signs of the elements of any circuit of $\mathcal{M}_{A}$ in the following way. It is well-known that the chirotope of a uniform oriented matroid relates the basis and the circuits as follows: $\chi(B)=-X_{j_{i}}\cdot X_{j_{i+1}}\cdot\chi(B^{\prime}),$ where $\underline{X}=\{j_{1},...,j_{r+1}\}$, $B=\underline{X}\setminus j_{i}$ and $B^{\prime}=\underline{X}\setminus j_{i+1}$ see [BVSWZ99, Section 3.5]. Thus, it can be seen that the signs of any circuit $\underline{X}=\{j_{1},...,j_{r+1}\}$ in ${\mathcal{M}}_{A}$ can be obtained from matrix $A$ by the equation
@@
see [R01, Lemma 2.1]. We also note the following.
@@
The Top and Bottom travels of a matrix $A$ were defined and studied in [R01] to detect when $\mathcal{M}_{A}$ is $1$-neighborly. Next, we will use them to determine when $\mathcal{M}_{A}$ is $k$-neighborly (Lemma 2.3).
@@
[R01, Lemma 3.1] There is a bijection between the set of all plain travels of the matrix $A$ and the set of all acyclic reorientations of $\mathcal{M}_{A}$.
@@
In this subsection we answer Conjecture 1.2 in the affirmative for LOMs when $r=2k+2\geq 4$ and $n\geq 3r-4$ (Theorem 2.11). The case $r=2k+1\geq 3$ and $n\geq r+2$ was proved for any oriented matroid in [HKM24, Corollay 3.17].
Let denote by $PT$ the set of plain travels of a matrix $A$ and let $PT^{k}$ be the set of plain travels of $A$ that are $k$-neighborly. More generally, given any set $\mathcal{P}$ of plain travels of $\mathcal{M}$, we denote by $\mathcal{P}^{k}$ the subset of plain travels of $\mathcal{P}$ that are $k$-neighborly. We notice that for a LOM $\mathcal{M}$, we have that $f_{\mathcal{M}}(r,n,k)=2|PT^{k}|$ by Lemma 2.5 and Remark 2.6. Thus, we may rewrite [HKM24, Corollay 3.17] now in terms of $k$-neighborly plain travels as follows.
@@ 2.5. $k$-Roudneff’s conjecture for low ranks
In this section we prove $k$-Roudneff’s Conjecture for LOMs for the cases $r=7$, $k=2$ and $n\geq 11$ (Corollary 2.26), for $r=8$, $k=2,3$ and $n=11,12$ (Theorem 2.27), for $r=9$, $k=2$ and $n=12$ and also for $r=9$, $k=3$ and $n\geq 12$ (Corollary 2.29). We recall that 1.2 was solved for LOMs when $k=1$ in [MR15].
In [HKM24] it was proved that 1.2 can be reduced to a finite number of cases as shown in the following theorem.
@@
[HKM24, Theorem 3.16] If $f_{\mathcal{M}^{\prime}}(r^{\prime},n^{\prime},k)\leq c_{r^{\prime}}(n^{\prime},k)$ for every rank $r^{\prime}\leq r$ uniform oriented matroid $\mathcal{M}^{\prime}$ on $n^{\prime}=2(r^{\prime}-k)+1$ elements, then
@@
[HKM24, Lemma 3.15] Let $\mathcal{M}$ be rank $r\geq 2k+1\geq 3$ uniform oriented matroid and let $e\in E$. Then,
@@
[BVSWZ99, Proposition 7.6.7] The family of LOMs is closed under contraction and deletion of elements.
@@
In [HKM24, Subsection 4.1], the authors provided a computer program to obtain $f_{\mathcal{M}}(r,n,k)$ from its chirotope, available at [supplemental_data] (in fact they obtained it via something called the $o$-vector whose $i$-th entry is $o_{i}$ with $i=0,\ldots,\lfloor\frac{r-1}{2}\rfloor$ and $o_{i}$ is the number of $i$-neighborly reorientations of $\mathcal{M}$ that are not $i+1$-neighborly). Now, for a LOM $\mathcal{M}_{A}$, we may use Equation (2) to obtain the signature of its basis an so its chirotope. Thus, using the computer program mentioned above we obtain $f_{\mathcal{M}_{A}}(r,n,k)$.
By Remark 1.3 it is enough to check Conjecture 1.2 only for the reorientation classes $\mathcal{M}$] of $\mathcal{M}$. So, in the case of LOMs it is sufficient by Theorem 2.18 to choose a representative matrix of a chessboard to obtain a representative matrix of its corresponding reorientation class. Hence, the computer program (available at [supplemental_data]) has to compute $2^{(n-r-1)(r-1)}$ different chirotopes to obtain $f_{\mathcal{M}_{A}}(r,n,k)$ for the case of LOMs by Theorem 2.20.
@@ 2.5.2. $k$-Roudneff’s conjecture for LOMs with $r=7$, $k=2$ and $n\geq 11$
We first notice that 1.2 was already proved for $r=7$, $k=2$ and $n=10$ in [HKM24, Theorem 4.1 (d)]. Now, in order to prove the conjecture for LOMs with $r=7$, $k=2$ and $n\geq 11$ it is enough to prove only the case $n=11$ by Proposition 2.24 since for $r<7$ and $k=2$ 1.2 was proved in [HKM24]. As mentioned above, we have to analyze $2^{(n-r-1)(r-1)}=2^{18}=262$ $144$ different matrices. We implemented our computer program and we resume the results in the following theorem.
@@
On the other hand, notice that 1.2 holds for $r=7$, $k=3$ since it was already proved for any oriented matroid in [HKM24, Theorem 3.17].
@@ 2.5.3. $k$-Roudneff’s conjecture for LOMs with $r=8$, $k=2,3$ and $n=11,12$
On the other hand, notice that 1.2 holds for $r=9,k=4$ since it was already proved for any oriented matroid in [HKM24, Theorem 3.17]. Finally, we observe that in order to prove the conjecture for LOMs with $r=9$ $k=2$ and $n\geq 15$ (using Proposition 2.24) the computer program has to calculate $2^{40}$ different matrices by Theorem 2.20.
@@ 3. $k$-Roudneff’s conjecture holds asymptotically
We will prove in Theorem 3.9 that $f(r,n,k)\leq F(r,n,k)$ for every $r\geq 2k+2\geq 4$ and $n\geq r+1\geq 5$ but first, we need some previous results. $k$-Roudneff’s conjecture was proved for $r=2k+1\geq 3$ and $n\geq r+2$ in [HKM24, Corollary 3.17]. On the other hand, the value of $c_{r}(n,k)$ was computed for $n\geq r+1$ when $r=2k+1$ in [HKM24, Theorem 3.11 and Proposition 3.12]. Combining these results and Remark 1.3 we obtain the following result that will be useful in this section.
@@
[HKM24] Let $\mathcal{M}$ be a rank $r=2k+1\geq 3$ oriented matroid on $n\geq r+1$ elements, then
@@
The value of $f(r,r,k)$ makes sense if we consider the oriented matroid as an arrangement. As we pointed out in the introduction (Subsection 1.2), an oriented matroid on $n$ elements of rank $r$ is naturally associated with some arrangement of $n$ (pseudo) hyperplanes in the projective space of dimension d=r1d=r-1. The arrangement decomposes the projective space into $d$-cells which in turn correspond to the acyclic reorientations (i.e., 0-neighborly reorientations) of the oriented matroid (We refer the reader to [FL78] to more details on the Topological Representation Theorem). In this sense, it is well-known that $f(r,r,0)\leq 2^{r-1}$ (see [Cor80]).
@@ Acknowledgements.
R. Hernández-Ortiz and L. P. Montejano were supported by SGR Grant 2021 00115 and by the project HERMES, funded by INCIBE and the European Union NextGenerationEU/PRTR. L. P Montejano was supported by the Ministry of Education and Science, Spain, under the project PID2023-146925OB-I00 and by the proyect ACoGe: PID2022-137283NB-C22.